Smithson Cutting is opening a new line of scissors for supermarket distribution. It estimates its fixed cost to be $ 450.00 and its variable cost to be $ 0.55 per unit. Selling price is expected to average $ 0.65 per unit. ​a) For Smithson​ Cutting, the​ break-even point in units​ = nothing units ​(enter your response as a whole​ number).
Answer #1

Total Cost = Fixed cost + no. units * variable cost

= 450 + X * 0.55

Revenue = no. of units * Selling price

= X * 0.65

Break even condition:

Revenue = total cost  

0.65X = 450 + 0.55X

X = 4500

hence break even point in units = 4500